OpenStudy (abmon98):

given that logax^2y=p and that loga(x/y^2)=q find logax and logay in terms of p and q and hence express logaxy in terms of p and q

OpenStudy (abmon98):

\[\log_{a} x^2y=p\] \[\log_{a} (x/y^2)=q\] \[\log_{a} x^2y=\log_{a} b^p\] \[x^2y=a^p\] \[\log_{a} (x/y^2)=\log_{a} a^q\] \[x/y^2=a^q\]